OpenStudy (anonymous):

pls help me...for my retrice.the sum of two numbers is 95, twice the larger increase by 3 times the smaller, give the sum of 232?

OpenStudy (amistre64):

for your what?

OpenStudy (anonymous):

the first equation is x + y=95.. now let x be larger and y be smaller, then the second equation is 2x+3y=232...solve simultaneously u get x=53,y=42

OpenStudy (amistre64):

l + s = 95 ; *-2 2l + 3s = 232 -2l - 2s = -190 2l +3s = 232 --------------- s = 42

OpenStudy (anonymous):

for my assignment i mean